∮|z|=1 (z+z̄)dz


1
Expert's answer
2021-08-19T15:51:02-0400

"\int _{\left|z\right|}^12z+z'\\
=\int _{\left|z\right|}^12z+0\\
=2z-2z\left|z\right|"